# FAB-PPI: Frequentist, Assisted by Bayes, Prediction-Powered Inference

### v1

@@ Acknowledgements
The alphafold dataset contains the following features for $N=10802$ protein residues analysed by Bludau et al. (2022): whether the residue is phosphorylated ($Z_{i}\in\{0,1\})$, whether the residue is part of an intrinsically disordered region (IDR, $Y_{i}\in\{0,1\}$), and the prediction of the AlphaFold model (Jumper et al., 2021) for the probability of $Y_{i}$ being equal to one ($f(X_{i})\in[0,1]$). The goal is to estimate the odds ratio of a protein being phosphorylated and being part of an IDR region, i.e.
@@
where $\mu_{1}=\mathrm{Pr}(Y=1\mid Z=1)$ and $\mu_{0}=\mathrm{Pr}(Y=1\mid Z=0)$. Following Angelopoulos et al. (2023b), given $\alpha\in(0,1)$, we construct $21-\alpha/$ 2 confidence intervals $\mathcal{C}_{0}=[l_{0},u_{0}]$ and $\mathcal{C}_{1}=[l_{1},u_{1}]$ for $\mu_{0}$ and $\mu_{1}$, respectively. Then, by a union bound, the interval
$\mathcal{C}=\left\{\frac{c_{1}}{1-c_{1}}\cdot\frac{1-c_{0}}{c_{0}}\colon c_{0}\in\mathcal{C}_{0},c_{1}\in\mathcal{C}_{1}\right\}=\left(\frac{l_{1}}{1-l_{1}}\cdot\frac{1-u_{0}}{u_{0}},\frac{u_{1}}{1-u_{1}}\cdot\frac{1-l_{0}}{l_{0}}\right)$
has coverage at least $1-\alpha$.
@@
The forest dataset contains the following features for $N=1596$ parcels of land in the Amazon rainforest examined during field visits (Bullock et al., 2020): whether the parcel has been subject to deforestation ($Y_{i}\in\{0,1\}$) and the prediction of a gradient-boosted tree model for the probability of $Y_{i}$ being equal to one ($f(X_{i})\in[0,1]$). The goal is to estimate the fraction of Amazon rainforest lost to deforestation, i.e. $\theta^{*}=\mathbb{E}[Y]$.
@@
The galaxies dataset contains the following features for $N=16743$ images from the Galaxy Zoo 2 initiative (Willett et al., 2013): whether the galaxy has spiral arms ($Y_{i}\in\{0,1\}$) and the prediction of a ResNet50 model (He et al., 2016) for the probability of $Y_{i}$ being equal to one ($f(X_{i})\in[0,1]$). The goal is to estimate the fraction of galaxies with spiral arms, i.e. $\theta^{*}=\mathbb{E}[Y]$.
@@
The genes dataset contains the following features for $N=61150$ gene promoter sequences: the expression level of the gene induced by the promoter and the prediction of a transformer model for the same quantity (Vaishnav et al., 2022). The goal is to estimate the median expression level across genes.
@@
The census dataset contains the following features for $N=380091$ individuals from the 2019 California census: the individual’s age, sex, and yearly income, as well as the prediction of a gradient-boosted tree model trained on the previous year’s raw data for the individual’s income. The goal is to estimate the ordinary least squares (OLS) regression coefficients when regressing income on age and sex.
@@
The healthcare dataset contains the following features for $N=318215$ individuals from the 2019 California census: the individual’s yearly income and whether they have health insurance ($Y_{i}\in\{0,1\}$), as well as the prediction of a gradient-boosted tree model trained on the previous year’s raw data for the probability of $Y_{i}$ being equal to one ($f(X_{i})\in[0,1]$). The goal is to estimate the logistic regression coefficient when regressing health insurance status on income.
@@
Compared to Figure 1, we include results for the non power-tuned methods, as well as for the ones that take into account the uncertainty on the measure of fit $m_{\theta}$ (i.e. PPI (full) and PPI++ (full)). In this example, power-tuning does not play a significant role and the same conclusions as in Section 5.1 hold. In particular, standard PPI induces shorter CIs than classical inference with constant volume across bias levels. On the other hand, FAB methods induce shorter CIs when the predictions are good. As the prediction bias increases, the volume of the FAB CIs with Gaussian prior grows unbounded, while the horseshoe prior eventually reverts to the PPI intervals. Furthermore, the coverage plot shows that the methods tested achieve similar coverage to the nominal level and to PPI (full) and PPI++ (full).
@@
Compared to Figure 2, we include results for the methods that use the Gaussian prior (FAB-PPI (N) and FAB-PPI++ (N)) and those that take into account the uncertainty on the measure of fit $m_{\theta}$ (i.e. PPI (full) and PPI++ (full)). Like the CI volume plots in the main text, the MSE plots clearly show the benefits of both power-tuning and adaptive shrinkage through the horseshoe prior: as $\sigma_{Y}$ increases, the power-tuned methods clearly outperform the standard alternatives, while shrinkage always helps compared to standard PPI because the predictions remain unbiased. In this case, the Gaussian prior performs similarly to the horseshoe as the prediction rule $f$ is unbiased. The coverage plots confirm that all methods achieve comparable coverage across noise levels.
@@
Figure S8: Full results for mean estimation experiment on real data. The left, middle, and right panels correspond to the alphafold, galaxies, and forest datasets, respectively. The top, middle, and bottom rows show average MSE, CI volume, and CI coverage, respectively over $1000$ repetitions for $\alpha=0.1$.
Compared to Figure 3, we include results for the non power-tuned methods, as well as for the ones that take into account the uncertainty on the measure of fit $m_{\theta}$ (i.e. PPI (full) and PPI++ (full)). The results are consistent with those presented in Section 5.2. In particular, FAB methods outperform the standard PPI alternatives and classical inference, while achieving comparable coverage. For the datasets and the values of $n$ considered, power-tuned methods perform similarly to the non-tuned ones. Among the FAB methods, the horseshoe and Gaussian priors achieve similar performance.
@@
Figure S9: Each subfigure includes 10 randomly chosen intervals for the classical, PPI++ and FAB-PPI++ methods. The left, middle, and right panels refer to the alphafold, galaxies, and forest datasets, respectively. The top, middle, and bottom rows refer correspond to different values of $n$.
@@
We repeat the mean estimation experiment on the forest dataset while varying the scale of the horseshoe prior used for FAB-PPI++ in Section S6.2.1. In addition to the scale $\widehat{\sigma}$ used in the main text, we consider the sample independent scale $1/\sqrt{n}$ and the data independent scale $1$. As already mentioned, the computation of the FAB-PPI confidence regions under a horseshoe prior with scale other than $\widehat{\sigma}$ involves numerical integration to compute the corresponding marginal likelihood. Figure S10 shows the average MSE, CI volume, and CI coverage for each of these choices, as well as for classical inference and PPI++.
@@
The predictions contained in this dataset are highly biased and this is reflected in the performance of the FAB-PPI methods. In particular, the Gaussian prior underperforms both classical inference and standard PPI, while the horseshoe prior achieves similar performance to standard PPI thanks to its robustness against large bias levels.
@@
More specifically, panel (a) and (b) corresponds to the OLS parameters associated with the age and sex covariates, respectively. On the one hand, FAB-PPI seems to perform well for the sex covariate, with similar performance between the Gaussian and horseshoe priors, and slightly improved MSE and CI volume compared to classical inference and standard PPI. On the other hand, the performance of FAB-PPI for the age covariate seems to be affected by bias in the dataset predictions. In particular, FAB-PPI under the Gaussian prior underperforms the alternatives for all $n$. On the other hand, while the horseshoe prior achieves worse performance than the other methods for small $n$, its performance improve as $n$ grows, and it eventually matches standard PPI. This suggests that, as $n$ increases and $\mathrm{var}(\widehat{\Delta}_{\theta})$ decreases, the observed value of the rectifier is increasingly considered as extreme, causing the influence from the horseshoe prior to eventually vanish thanks to its robustness to extreme bias levels.

### v2

@@ Acknowledgements
The alphafold dataset contains the following features for N=10 802N=10\,802 protein residues analysed by Bludau et al. (2022): whether the residue is phosphorylated ($Z_{i}\in\{0,1\}$), whether the residue is part of an intrinsically disordered region (IDR, $Y_{i}\in\{0,1\}$), and the prediction of the AlphaFold model (Jumper et al., 2021) for the probability of $Y_{i}$ being equal to one ($f(X_{i})\in[0,1]$). The goal is to estimate the odds ratio of a protein being phosphorylated and being part of an IDR, i.e.
@@
where $\mu_{1}=\mathrm{Pr}(Y=1\mid Z=1)$ and $\mu_{0}=\mathrm{Pr}(Y=1\mid Z=0)$. Following Angelopoulos et al. (2023a), given $\alpha\in(0,1)$, we construct $1-\alpha/2$ confidence intervals $\mathcal{C}_{0}=[l_{0},u_{0}]$ and $\mathcal{C}_{1}=[l_{1},u_{1}]$ for $\mu_{0}$ and $\mu_{1}$, respectively. Then, by a union bound, the interval
$\mathcal{C}=\left\{\frac{c_{1}}{1-c_{1}}\cdot\frac{1-c_{0}}{c_{0}}\colon c_{0}\in\mathcal{C}_{0},c_{1}\in\mathcal{C}_{1}\right\}=\left[\frac{l_{1}}{1-l_{1}}\cdot\frac{1-u_{0}}{u_{0}},\frac{u_{1}}{1-u_{1}}\cdot\frac{1-l_{0}}{l_{0}}\right]$
has coverage at least $1-\alpha$. Note that the union bound above may result in a conservative confidence interval, leading to coverage significantly larger than $1-\alpha$ in practice, as in the left panel of Figure 3.
@@
The forest dataset contains the following features for N=1 596N=1\,596 parcels of land in the Amazon rainforest examined during field visits (Bullock et al., 2020): whether the parcel has been subject to deforestation ($Y_{i}\in\{0,1\}$) and the prediction of a gradient-boosted tree model for the probability of $Y_{i}$ being equal to one ($f(X_{i})\in[0,1]$). The goal is to estimate the fraction of Amazon rainforest lost to deforestation, i.e. $\theta^{\star}=\mathbb{E}[Y]$.
@@
The galaxies dataset contains the following features for N=16 743N=16\,743 images from the Galaxy Zoo 2 initiative (Willett et al., 2013): whether the galaxy has spiral arms ($Y_{i}\in\{0,1\}$) and the prediction of a ResNet50 model (He et al., 2016) for the probability of $Y_{i}$ being equal to one ($f(X_{i})\in[0,1]$). The goal is to estimate the fraction of galaxies with spiral arms, i.e. $\theta^{\star}=\mathbb{E}[Y]$.
@@
The genes dataset contains the following features for N=61 150N=61\,150 gene promoter sequences: the expression level of the gene induced by the promoter and the prediction of a transformer model for the same quantity (Vaishnav et al., 2022). The goal is to estimate the median expression level across genes.
@@
The census dataset contains the following features for N=380 091N=380\,091 individuals from the 2019 California census: the individual’s age, sex, and yearly income, as well as the prediction of a gradient-boosted tree model trained on the previous year’s raw data for the individual’s income. The goal is to estimate the ordinary least squares (OLS) regression coefficients when regressing income on age and sex.
@@
The healthcare dataset contains the following features for N=318 215N=318\,215 individuals from the 2019 California census: the individual’s yearly income and whether they have health insurance ($Y_{i}\in\{0,1\}$), as well as the prediction of a gradient-boosted tree model trained on the previous year’s raw data for the probability of $Y_{i}$ being equal to one ($f(X_{i})\in[0,1]$). The goal is to estimate the logistic regression coefficient when regressing health insurance status on income.
S5.2 Implementation
Code implementing the FAB-PPI method is written in Python and made available at https://github.com/stefanocortinovis/fab-ppi. Comparisons with standard PPI are performed using the ppi-py package (Angelopoulos et al., 2023b). All of the experiments presented here were run locally on an Intel Core i7-11850H CPU.
@@
Compared to Figure 1, we include results for the non power-tuned methods, as well as for the ones that take into account the uncertainty in the measure of fit $m_{\theta}$ (i.e. PPI (full) and PPI++ (full)). In this example, power tuning does not play a significant role and the same conclusions as in Section 5.1 hold. In particular, standard PPI induces shorter CIs than classical inference with constant volume across bias levels. On the other hand, FAB methods induce shorter CIs when the predictions are good. As the prediction bias increases, the volume of the FAB CIs with Gaussian prior grows without bound, while the horseshoe prior eventually reverts to the PPI intervals. Furthermore, the coverage plot shows that the methods tested achieve similar coverage to the nominal level and to PPI (full) and PPI++ (full).
@@
Compared to Figure 2, we include results for the methods that use the Gaussian prior (FAB-PPI (N) and FAB-PPI++ (N)) and those that take into account the uncertainty in the measure of fit $m_{\theta}$ (i.e. PPI (full) and PPI++ (full)). Like the CI volume plots in the main text, the MSE plots clearly show the benefits of both power tuning and adaptive shrinkage through the horseshoe prior: as $\sigma_{Y}$ increases, the power-tuned methods clearly outperform the standard alternatives, while shrinkage always helps compared to standard PPI because the predictions remain unbiased. In this case, the Gaussian prior performs similarly to the horseshoe as the prediction rule $f$ is unbiased. The coverage plots confirm that all methods achieve comparable coverage across noise levels.
@@
Figure S8: Full results for mean estimation experiment on real data. The left, middle, and right panels correspond to the alphafold, galaxies, and forest datasets, respectively. The top, middle, and bottom rows show average MSE, CI volume, and CI coverage, respectively, over 1 0001\,000 repetitions for $\alpha=0.1$.
Compared to Figure 3, we include results for the non power-tuned methods, as well as for the ones that take into account the uncertainty in the measure of fit $m_{\theta}$ (i.e. PPI (full) and PPI++ (full)). The results are consistent with those presented in Section 5.2. In particular, FAB methods outperform the standard PPI alternatives and classical inference, while achieving comparable coverage. For the datasets and the values of $n$ considered, power-tuned methods perform similarly to the non-tuned ones. Among the FAB methods, the horseshoe and Gaussian priors achieve similar performance.
@@
Figure S9: Each subfigure includes 10 randomly chosen intervals for the classical, PPI++ and FAB-PPI++ methods. The left, middle, and right panels refer to the alphafold, galaxies, and forest datasets, respectively. The top, middle, and bottom rows correspond to different values of $n$.
@@
We repeat the mean estimation experiment on the forest dataset while varying the scale of the horseshoe prior used for FAB-PPI++ in Section S6.2.1. In addition to the scale $\widehat{\sigma}$ used in the main text, we consider the sample-independent scale $1/\sqrt{n}$ and the data-independent scale $1$. As already mentioned, the computation of the FAB-PPI confidence regions under a horseshoe prior with scale other than $\widehat{\sigma}$ involves numerical integration to compute the corresponding marginal likelihood. Figure S10 shows the average MSE, CI volume, and CI coverage for each of these choices, as well as for classical inference and PPI++.
@@
The predictions contained in this dataset are highly biased, and this is reflected in the performance of the FAB-PPI methods. In particular, the Gaussian prior underperforms both classical inference and standard PPI, while the horseshoe prior achieves similar performance to standard PPI thanks to its robustness against large bias levels.
@@
More specifically, panels (a) and (b) correspond to the OLS parameters associated with the age and sex covariates, respectively. On the one hand, FAB-PPI seems to perform well for the sex covariate, with similar performance between the Gaussian and horseshoe priors, and slightly improved MSE and CI volume compared to classical inference and standard PPI. On the other hand, the performance of FAB-PPI for the age covariate seems to be affected by bias in the dataset predictions. In particular, FAB-PPI under the Gaussian prior underperforms the alternatives for all $n$. On the other hand, while the horseshoe prior achieves worse performance than the other methods for small $n$, its performance improves as $n$ grows, and it eventually matches standard PPI. This suggests that, as $n$ increases and $\mathrm{var}(\widehat{\Delta}_{\theta})$ decreases, the observed value of the rectifier is increasingly considered as extreme, causing the influence from the horseshoe prior to eventually vanish thanks to its robustness to extreme bias levels.